# InstructPipe: Building Visual Programming Pipelines with Human Instructions Using LLMs

### v2

@@ 4.2. Node Selector
Node Selector filters out unrelated nodes by providing the LLM with a short description of each node. Fig. 5 shows the prompt we use in Node Selector. Followed by a general task description and guidelines, we list all node types with a short description that explains the function of each. Several nodes come with recommendation(s) when the users interact with Visual Blocks, and we also include such node recommendations in the prompt. The main intuition of this prompt design is based on how existing open-source libraries (e.g., Numpy (Harris et al., 2020)) present a high-level overview of all functions [^2] See an example in the following link: https://numpy.org/doc/1.25/reference/routines.array-manipulation.html. The documentation typically presents a list of supported functions (in each category), followed by a short description so that developers can quickly find their desired functions. At the end of the prompt, we provide a list of Q&As as few-shot examples to support the LLM to learn and adapt to the context of the task.
@@ 4.3. Code Writer
With a pool of selected nodes, the Code Writer module can write pipeline rendering pseudocode. Fig. 6 shows the structure of the prompt utilized in this LLM module. Similar to section 4.2, the prompt starts with a general introduction and several guidelines. The major difference in the prompt design in this stage lies in the granularity of each node introduction. We provide a detailed configuration for each selected node with additional information, including 1) input data types, 2) output data types, and 3) an example, represented in pseudocode, for how this node connects to other nodes. We include a detailed explanation of the full node configuration in Section B.1.2. Similar to the previous LLM module (section 4.2), the prompt design here is also inspired by the documentation of existing software libraries. Specifically, we gain inspiration from low-level function-specific documentation [^3] See an example in numpy.shape: https://numpy.org/doc/1.25/reference/generated/numpy.shape.html#numpy-shape, which typically includes 1) a detailed description, 2) data types in the input/output, followed by 3) one or more examples of a few lines of code for how developers can use the function.
The prompt also includes a Q&A list as few-shot examples. However, providing few-shot examples in this stage is non-trivial. The reason lies in the dynamics of the node selection pool. A combination of all the nodes causes many possible options, and it is impossible to design a dedicated list of few-shot examples in each possible case. Therefore, we only created an example list for each pipeline tag (i.e., “language”, “visual”, and “multimodal”) and intended to utilize these few-shot pipelines to teach LLMs example use cases in each category. This implies that in-context pipelines may include nodes that were not selected for the prompt. This can potentially lead to LLM hallucinations (Huang et al., 2023b), i.e., utilizing the nodes that do not exist in our node library. We mitigated this issue by adding specific prompts that explicitly show a list of supported nodes (i.e., the contents start with “the following is a full list of …” in Fig. 6). However, LLM hallucination is a community-wise challenge, and we also find that our approach cannot completely eliminate this issue in visual programming. Therefore, InstructPipe conduct a sanity check for the Code Writer outputs and directly disposes of the line of pseudocode with such hallucinated nodes. This can ensure that the generated code is in a valid data format for rendering the pipeline in Visual Blocks.
@@ 4.4. Code Interpreter
    Graph Generation with Default Node Parameters: We generated a DAG based on the tokenized results, and apply predefined default node parameters in each generated node. For example, by default, the temperature and the max output tokens for the PaLM node are set to 0.5 and 256, respectively. If users are not satisfied with the default values, they can interactively adjust the parameters in the node-graph editor.
@@ 5. Technical Evaluation
InstructPipe contributes a framework for generating specifications for visual programming pipelines based on text prompts from users. To characterize the system’s performance, we designed a technical evaluation to asses the accuracy of the generated graphs for a variety of prompts.
@@ 5.1. Data Collection
To compute the accuracy of our generated pipelines, we need to collect a corpus with pairs of instructions and their corresponding ground-truth pipelines. Therefore, we organized a two-day hybrid workshop with 23 participants, aiming to collect real pipelines that Visual Blocks users would build for their creative usage. The 23 participants (F: 6; M: 17) are composed of five software engineers, four research scientists, four students, three designers, two project managers, and two engineering mangers. Six attendees claimed that they had prior experience in using Visual Blocks. As this was a data collection study rather than a user study, where each participant here served as a data creator and annotator, we did not restrict participation to individuals who self-identified as novices. The workshop began with a 15-minute Visual Blocks tutorial walking the participants through the nodes and the pipeline-building process. After the tutorial, attendees created pipelines independently. Once they finished creating the pipelines, participants were required to caption their pipelines and upload them. We utilized this corpus of data pairs (caption/pipeline) as the data set for the technical evaluation.
The workshop was an open-ended creation process, in which participants were free to use any node available in Visual Blocks with more than the 27 nodes covered by InstructPipe. The InstructPipe feature was not available in the workshop. After the workshop, we post-processed our collected data and achieved 48 pipelines (23 language pipelines, seven visual pipelines and 18 multi-modal pipelines) for our technical evaluations. The post-processing procedure details are presented in Appendix C.1.
@@ 5.2. Metric: The Number of User Interactions
To quantify the efficacy of InstructPipe based on our goal of accelerating and streamlining pipeline creation, we defined the metric Number of User Interactions as follow:
@@ 5.3. Experiment Setups and Results
We ran our generation algorithm on the pipeline captions six times (three times for each caption $\times$ two captions for each pipeline), and computed an averaged performance among the six trials for each pipeline. Table 1 summarizes the results of the technical evaluation. Compared to building a pipeline from scratch, InstructPipe allows the user to complete a pipeline with 18.9% of the user interactions, demonstrating the potential of InstructPipe to require more than 5X fewer interactions. Seven generated pipelines directly satisfied with instructions without user interactions in all six trials, and 38 generated pipelines completed at least once in any of the six trials.
@@ 6. User Evaluation
While the technical evaluation demonstrates the accuracy of InstructPipe among various real pipelines created by participants, it is still unclear what is the actual user experience when real users go through the entire system workflow. Therefore, we conducted an in-person user study of InstructPipe with another group of participants, aiming to provide more insights into our system performance as well as explore new user experiences brought by InstructPipe. The study recruitment was in accordance with the ethics board of our institution. We obtained participant consent before the study began.
@@ 6.1. Study Design
In the user evaluation, we aimed to investigate how the interface condition (with InstructPipe and without InstructPipe; the independent variable) affects the user experience and behaviors (the dependent variable). We will refer to these two interface conditions as “InstructPipe” and “Visual Blocks” in the following content. Fig. 7 visualizes the complete study flow. In each condition, participants completed the two pipelines with counterbalance (referred to as Task 1 and Task 2 in Fig. 7).
@@
We carefully designed the experiment to create a fair study that could be completed with reasonable effort. In the following content, we elaborate how we made two important decisions related to the study’s rigor:
@@ 6.1.1. Two controlled pipelines with full counterbalancing
Our user evaluation focuses on two controlled pipelines with full counterbalancing. While we acknowledge that more pipelines (e.g., four, six, or more) could enhance generalizability, such designs would also inevitably increase the size of the required user groups, even without fully counterbalancing. For example, fully counterbalancing four controlled pipelines require 12×12$\times$12 × more participants. Partially counterbalancing with four pipelines using the Latin Square design still requires us to double the number of participants. Additionally, novice participants are likely to progressively gain experiences within the study, and such learning effects will weaken the design of partial counterbalancing. We believe that two pipelines with fully counterbalancing is a reasonable experiment setup in this work, and future work could consider extending and scaling up these experiments.
@@ 6.2. Procedure
Each study session takes 55 - 65 minutes in total. The study started with 10-15 minute hands-on training for both conditions. The training included 1) all the Visual Blocks interactions needed to complete the subsequent steps of the experiment, and 2) all the nodes that participants will need to use for pipeline creation in the main session. Participants were also encouraged to experiment with building a pipeline independently, and to ask questions.
@@
During the task, participants were allowed to consult with us for technical help. If participants were unable to make progress, we provided hints. We provided much more hints in the baseline condition, and we made this decision to ensure every novice-level participant can finish their tasks within a reasonable amount of time. Appendix D.3 contains more details and discussion of the assistance we provided in the study. As an optional extension to the study, eight participants were offered an open-ended pipeline creation, where participants prototyped their own ideas using InstructPipe. This optional section was offered based on the progress of the participant in the previous sections, and time constraints so that the study duration was controlled within the time we guaranteed in our recruitment process.
After conducting all pipeline-condition combinations, participants answered open-ended questions in a semi-structured interview. The interview script is available in the appendix D.1. Participants provided their general impression for each condition, listed pros and cons, identified potential future use cases, and critiqued the user interface for future improvements. We transcribed the interviews and conducted the open coding analysis on the qualitative data (Strauss, 1987; Strauss et al., 1990). More specifically, we categorized the quotes based on our observations and then refined the code for presentation.
@@ 6.3. Participants
We recruited 16 participants from our internal participant pool, which is specifically designed for UX research within our institution. Importantly, none of the participants was involved in our project, and authors in charge of the study did not personally know any of the participants. We screened participants on their self-reported programming experience and machine learning skills. All of the 16 selected participants rated their “Programming Experience” and “’Machine Learning Skill’ as “Intermediate” or below (See Table 3 for a full breakdown). We intentionally recruited novice users, as we envision them as intended users of InstructPipe.
@@ 6.4.1. Task Completion Time
Back-end logs were used to collect timestamps for starting and ending events. Then, completion time for each condition were calculated per task for each participant.
@@ 6.5.1. InstructPipe Reduces Users’ Workload
Fig. 8 further visualizes the results of users’ perceived workload in six sub-scales. The Wilcoxon signed ranks test revealed significant differences on five sub-scales, all except “Mental Demand” (see section 7.2 for more explanations and discussion). Furthermore, the test indicates that all participants unanimously considered that InstructPipe provides lower or equal workload on the subscales of “Physical Demand”, “Temporal Demand”, “Performance” and “Effort” ($W=0$). These quantitative results, with both objective and subjective metrics, demonstrate the potential for InstructPipe to dramatically reduce users’ workload during the pipeline creation process.
@@ 6.5.2. On-boarding Support of Visual Programming
P1’s comment matches with our observation of participants’ behaviors during the study. In the Visual Blocks condition, we observed that people were more easily stuck in their creative purposes, which required our support. Typical support included 1) guiding participants if they went too far away from the correct pipeline, and 2) reminding them of an important node for the pipeline, although we introduced all the necessary nodes in our training session.
@@ 7.1. Human-AI Collaboration in Prototyping Open-ended ML Pipelines
Such takeaways differ from early-stage findings that show LLMs can achieve full automation of ML inference (Surís et al., 2023; Gupta and Kembhavi, 2023). The main reason is that existing work built their ad hoc solutions for target use scenarios, respectively. In contrast, InstructPipe covers a larger range of ML models (section 3.2) and aims for an open-ended use case. Our results show that LLMs (we used GPT-3.5-turbo in the study) still fail to write robust code with prompt engineering techniques. This aligns with the latest research findings that shows that even the latest LLMs still fail to formulate a whole working pipeline (Wen et al., 2024; Qin et al., 2023b).
@@ 7.2.1. Instruction
P15’s comment in section 6.5.5 summarizes the first aspect that causes mental burden. Although the “instruction-to-pipeline” process is fast and seems effortless, the process of framing a prompt is one factor that may overwhelm users, especially those who are more accustomed to visual thinking. InstructPipe requires its users to 1) be clear about the problem they want to solve, and 2) be able to clearly articulate the desired pipeline. Such requirements cause a mental burden to the user. We believe that our results can reinforce Zamfirescu-Pereira et al.’s findings on how non-experts may not prompt LLMs well (Zamfirescu-Pereira et al., 2023) in the visual programming domain.
@@ 7.2.2. Perception
The integration of LLM support into the visual programming interface enables a “multimodal programming” experience (Dietz et al., 2023), in which, users can program through both verbal and visual approaches. However, this flexibility increases perceptual burden as users switch between visual and text-based thinking (Paivio, 1991). Interestingly, our results seemingly contradict psychological findings based on the Dual Coding Theory (DCT) that show a combination of verbal and visual information actually helps humans’ memory process [^4] For example, people feel it easier to remember a new word if they learn the word using a vocabulary card with a figure that explains the texts. (Paivio, 1969; Paivio et al., 2006). Therefore, we believe that the mental workload stems not from dislike of multimodal workspaces, but from the lack of a transparent interface that aligns users’ mental models with AI reasoning both verbally and visually. That being said, a next-generation copilot should visualize a pipeline (i.e., visual info) while the user is prompting the system (i.e., verbal info), constituting and interfacing multimodal processing in humans’ brain.
@@ 7.2.3. Debugging
When a rendered pipeline does not match users’ expectations, users need to debug (see P13’s comment section 6.5.5). Specifically, users need to “invest extra effort to review and understand the generated content” (Zhang et al., 2023a) and then solve the issues caused by the AI assistant. In essence, debugging is a professional programming skill, which understandably can be mentally overwhelming for beginner-level users. While InstructPipe visualizes generated code in the visual programming platform, our results suggests that future systems should provide more guidance for beginners to better proceed with their programming tasks.

### v3

@@ 4.2. Node Selector
Node Selector filters out unrelated nodes by providing the LLM with a short description of each node. Figure 5 shows the prompt we use in Node Selector. Followed by a general task description and guidelines, we list all node types with a short description that explains the function of each. Several nodes come with recommendation(s) when the users interact with Visual Blocks, and we also include such node recommendations in the prompt. The main intuition of this prompt design is based on how existing open-source libraries (e.g., Numpy (Harris et al., 2020)) present a high-level overview of all functions [^2] See an example in the following link: https://numpy.org/doc/1.25/reference/routines.array-manipulation.html. The documentation typically presents a list of supported functions (in each category), followed by a short description so that developers can quickly find their desired functions. At the end of the prompt, we provide a list of Q&As as few-shot examples to support the LLM to learn and adapt to the context of the task.
@@ 4.3. Code Writer
With a pool of selected nodes, the Code Writer module can write pipeline rendering pseudocode. Figure 6 shows the structure of the prompt utilized in this LLM module. Similar to section 4.2, the prompt starts with a general introduction and several guidelines. The major difference in the prompt design in this stage lies in the granularity of each node introduction. We provide a detailed configuration for each selected node with additional information, including 1) input data types, 2) output data types, and 3) an example, represented in pseudocode, for how this node connects to other nodes. We include a detailed explanation of the full node configuration in Section B.1.2. Similar to the previous LLM module (section 4.2), the prompt design here is also inspired by the documentation of existing software libraries. Specifically, we gain inspiration from low-level function-specific documentation [^3] See an example in numpy.shape: https://numpy.org/doc/1.25/reference/generated/numpy.shape.html#numpy-shape, which typically includes 1) a detailed description, 2) data types in the input/output, followed by 3) one or more examples of a few lines of code for how developers can use the function.
The prompt also includes a Q&A list as few-shot examples. However, providing few-shot examples in this stage is non-trivial. The reason lies in the dynamics of the node selection pool. A combination of all the nodes causes many possible options, and it is impossible to design a dedicated list of few-shot examples in each possible case. Therefore, we only created an example list for each pipeline tag (i.e., “language”, “visual”, and “multimodal”) and intended to utilize these few-shot pipelines to teach LLMs example use cases in each category. This implies that in-context pipelines may include nodes that were not selected for the prompt. This can potentially lead to LLM hallucinations (Huang et al., 2023b), i.e., utilizing the nodes that do not exist in our node library. We mitigated this issue by adding specific prompts that explicitly show a list of supported nodes (i.e., the contents start with “the following is a full list of …” in Figure 6). However, LLM hallucination is a community-wide challenge, and we also find that our approach cannot eliminate this issue in visual programming. Therefore, InstructPipe conducts a sanity check for the Code Writer outputs and directly disposes of the line of pseudocode with such hallucinated nodes. This can ensure that the generated code is in a valid data format for rendering the pipeline in Visual Blocks.
@@ 4.4. Code Interpreter
    Graph Generation with Default Node Parameters: We generated a DAG based on the tokenized results and applied predefined default node parameters in each generated node.
    For example, by default, the temperature and the max output tokens for the PaLM node are set to 0.5 and 256, respectively. If users are not satisfied with the default values, they can interactively adjust the parameters in the node-graph editor.
@@ 5. Technical Evaluation
InstructPipe contributes a framework for generating specifications for visual programming pipelines based on text prompts from users. To characterize the system’s performance, we designed a technical evaluation to assess the accuracy of the generated graphs for a variety of prompts.
@@ 5.1. Data Collection
To compute the accuracy of our generated pipelines, we need to collect a corpus with pairs of instructions and their corresponding ground-truth pipelines. Therefore, we organized a two-day hybrid workshop with 23 participants, aiming to collect real pipelines that Visual Blocks users would build for their creative usage. The 23 participants (F: 6; M: 17) are composed of five software engineers, four research scientists, four students, three designers, two project managers, and two engineering managers. Six attendees claimed that they had prior experience in using Visual Blocks. As this was a data collection study rather than a user study, where each participant here served as a data creator and annotator, we did not restrict participation to individuals who self-identified as novices. The workshop began with a 15-minute Visual Blocks tutorial walking the participants through the nodes and the pipeline-building process. After the tutorial, attendees created pipelines independently. Once they finished creating the pipelines, participants were required to caption their pipelines and upload them. We utilized this corpus of data pairs (caption/pipeline) as the data set for the technical evaluation.
The workshop was an open-ended creation process in which participants were free to use any node available in Visual Blocks with more than the 27 nodes covered by InstructPipe. The InstructPipe feature was not available in the workshop. After the workshop, we post-processed our collected data and achieved 48 pipelines (23 language pipelines, seven visual pipelines and 18 multi-modal pipelines) for our technical evaluations. The post-processing procedure details are presented in Appendix C.1.
@@ 5.2. Metric: The Number of User Interactions
To quantify the efficacy of InstructPipe based on our goal of accelerating and streamlining pipeline creation, we defined the metric Number of User Interactions as follows:
@@ 5.3. Experiment Setups and Results
We ran our generation algorithm on the pipeline captions six times (three times for each caption $\times$ two captions for each pipeline), and computed an averaged performance among the six trials for each pipeline.
Table 1 summarizes the results of the technical evaluation. Compared to building a pipeline from scratch, InstructPipe allows the user to complete a pipeline with 18.9% of the user interactions, demonstrating the potential of InstructPipe to require more than 5X fewer interactions. Seven generated pipelines directly satisfied with instructions without user interactions in all six trials, and 38 generated pipelines completed at least once in any of the six trials.
@@ 6. User Evaluation
While the technical evaluation demonstrates the accuracy of InstructPipe among various real pipelines created by participants, it is still unclear what is the actual user experience when real users go through the entire system workflow. Therefore, we conducted an in-person user study of InstructPipe with another group of participants, aiming to provide more insights into our system performance as well as explore new user experiences brought by InstructPipe. The study recruitment was in accordance with the ethics board of Google. We obtained participant consent before the study began.
@@ 6.1. Study Design
In the user evaluation, we aimed to investigate how the interface condition (with InstructPipe and without InstructPipe; the independent variable) affects the user experience and behaviors (the dependent variable). We will refer to these two interface conditions as “InstructPipe” and “Visual Blocks” in the following content. Figure 7 visualizes the complete study flow. In each condition, participants completed the two pipelines with counterbalance (referred to as Task 1 and Task 2 in Figure 7).
@@
We carefully designed the experiment to create a fair study that could be completed with reasonable effort. In the following content, we elaborate on how we made two important decisions related to the study’s rigor:
@@ 6.1.1. Two controlled pipelines with full counterbalancing
Our user evaluation focuses on two controlled pipelines with full counterbalancing. While we acknowledge that more pipelines (e.g., four, six, or more) could enhance generalizability, such designs would also inevitably increase the size of the required user groups, even without fully counterbalancing. For example, fully counterbalancing four controlled pipelines requires 12×12$\times$12 × more participants. Partially counterbalancing with four pipelines using the Latin Square design still requires us to double the number of participants. Additionally, novice participants are likely to progressively gain experiences within the study, and such learning effects will weaken the design of partial counterbalancing. We believe that two pipelines with full counterbalancing are a reasonable experiment setup in this work, and future work could consider extending and scaling up these experiments.
@@ 6.2. Procedure
Each study session takes 55 - 65 minutes in total. The study started with 10-15 minutes of hands-on training for both conditions. The training included 1) all the Visual Blocks interactions needed to complete the subsequent steps of the experiment, and 2) all the nodes that participants will need to use for pipeline creation in the main session. Participants were also encouraged to experiment with building a pipeline independently, and to ask questions.
@@
During the task, participants were allowed to consult with us for technical help. If participants were unable to make progress, we provided hints. We provided many more hints in the baseline condition, and we made this decision to ensure every novice-level participant can finish their tasks within a reasonable amount of time. Appendix D.3 contains more details and discussions of the assistance we provided in the study. As an optional extension to the study, eight participants were offered an open-ended pipeline creation, where participants prototyped their own ideas using InstructPipe. This optional section was offered based on the progress of the participant in the previous sections, and time constraints so that the study duration was controlled within the time we guaranteed in our recruitment process.
After conducting all pipeline-condition combinations, participants answered open-ended questions in a semi-structured interview. The interview script is available in the appendix D.1. Participants provided their general impression of each condition, listed pros and cons, identified potential future use cases, and critiqued the user interface for future improvements. We transcribed the interviews and conducted the open coding analysis on the qualitative data (Strauss, 1987; Strauss et al., 1990). More specifically, we categorized the quotes based on our observations and then refined the code for presentation.
@@ 6.3. Participants
We recruited 16 participants from our internal participant pool, which is specifically designed for UX research within our institution. Importantly, none of the participants was involved in our project, and the authors in charge of the study did not personally know any of the participants. We screened participants on their self-reported programming experience and machine learning skills. All of the 16 selected participants rated their “Programming Experience” and “’Machine Learning Skill’ as “Intermediate” or below (See Table 3 for a full breakdown). We intentionally recruited novice users, as we envision them as intended users of InstructPipe.
@@ 6.4.1. Task Completion Time
Back-end logs were used to collect timestamps for starting and ending events. Then, the completion time for each condition was calculated per task for each participant.
@@ 6.5.1. InstructPipe Reduces Users’ Workload
Figure 8 further visualizes the results of users’ perceived workload in six sub-scales. The Wilcoxon signed ranks test revealed significant differences on five sub-scales, all except “Mental Demand” (see section 7.2 for more explanations and discussion). Furthermore, the test indicates that all participants unanimously considered that InstructPipe provides lower or equal workload on the subscales of “Physical Demand”, “Temporal Demand”, “Performance” and “Effort” ($W=0$). These quantitative results, with both objective and subjective metrics, demonstrate the potential for InstructPipe to dramatically reduce users’ workload during the pipeline creation process.
@@ 6.5.2. On-boarding Support of Visual Programming
P1’s comment matches our observation of participants’ behaviors during the study. In the Visual Blocks condition, we observed that people were more easily stuck in their creative purposes, which required our support. Typical support included 1) guiding participants if they went too far away from the correct pipeline, and 2) reminding them of an important node for the pipeline, although we introduced all the necessary nodes in our training session.
@@ 7.1. Human-AI Collaboration in Prototyping Open-ended ML Pipelines
Such takeaways differ from early-stage findings that show LLMs can achieve full automation of ML inference (Surís et al., 2023; Gupta and Kembhavi, 2023). The main reason is that existing work built their ad hoc solutions for target use scenarios, respectively. In contrast, InstructPipe covers a larger range of ML models (section 3.2) and aims for an open-ended use case. Our results show that LLMs (we used GPT-3.5-turbo in the study) still fail to write robust code with prompt engineering techniques. This aligns with the latest research findings that show that even the latest LLMs still fail to formulate a whole working pipeline (Wen et al., 2024; Qin et al., 2023b).
@@ 7.2.1. Instruction
P15’s comment in section 6.5.5 summarizes the first aspect that causes mental burden. Although the “instruction-to-pipeline” process is fast and seems effortless, the process of framing a prompt is one factor that may overwhelm users, especially those who are more accustomed to visual thinking. InstructPipe requires its users to 1) be clear about the problem they want to solve, and 2) be able to clearly articulate the desired pipeline. Such requirements cause a mental burden to the user (Blackwell, 1996). We believe that our results can reinforce the existing knowledge on how non-experts may not prompt LLMs well (Zamfirescu-Pereira et al., 2023; Nguyen et al., 2024) in the visual programming domain.
@@ 7.2.2. Perception
The integration of LLM support into the visual programming interface enables a “multimodal programming” experience (Dietz et al., 2023), in which, users can program through both verbal and visual approaches. However, this flexibility increases perceptual burden as users switch between visual and text-based thinking (Paivio, 1991). Interestingly, our results seemingly contradict psychological findings based on the Dual Coding Theory (DCT) that show a combination of verbal and visual information actually helps humans’ memory process [^4] For example, people feel it easier to remember a new word if they learn the word using a vocabulary card with a figure that explains the texts. (Paivio, 1969; Paivio et al., 2006). Therefore, we believe that the mental workload stems not from dislike of multimodal workspaces, but from the lack of a transparent interface that aligns users’ mental models with AI reasoning both verbally and visually. That being said, a next-generation copilot should visualize a pipeline (i.e., visual info) while the user is prompting the system (i.e., verbal info), constituting and interfacing multimodal processing in humans’ brains.
@@ 7.2.3. Debugging
When a rendered pipeline does not match users’ expectations, users need to debug (see P13’s comment section 6.5.5). Specifically, users need to “invest extra effort to review and understand the generated content” (Zhang et al., 2023a) and then solve the issues caused by the AI assistant. In essence, debugging is a professional programming skill, which understandably can be mentally overwhelming for beginner-level users. While InstructPipe visualizes generated code in the visual programming platform, our results suggest that future systems should provide more guidance for beginners to better proceed with their programming tasks.